OpenStudy (anonymous):

Decide which part of the quadratic formula tells you whether the quadratic equation can be solved by factoring. −b b2 − 4ac 2a Use the part of the quadratic formula that you chose above and find its value, given the following quadratic equation: 2x2 + 7x + 3 = 0

pooja195 (pooja195):

\(b^2-4ac\) is called 'Discriminant' which decides that whether the the quadratic equation can be solved by factoring.

OpenStudy (anonymous):

I know that

pooja195 (pooja195):

So you need help factoring?

OpenStudy (anonymous):

Basically

pooja195 (pooja195):

Well what numbers multiply to make 6 but add to make 7?

OpenStudy (anonymous):

6 and 1

OpenStudy (anwaarullah):

2x^+6x+x+3=0

pooja195 (pooja195):

ok now you have (2x+6 )(2x+1) Can you divide both of these by a GCF?

pooja195 (pooja195):

What number can we divide 2 and 6 by?

OpenStudy (anonymous):

2

OpenStudy (anwaarullah):

take common

pooja195 (pooja195):

You cant simpyfy anything more so you have 2x+3=0 2x+1=0 solve both for x

OpenStudy (anwaarullah):

 solve more